Shift report 20030217 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Feb 17 2003 - 08:05:25 EST

  • Next message: Apache: "Shift report 20030217 08:00-16:00"
    ------------------------------------------------------------
    Date: Mon Feb 17 08:05:24 2003, Shift: 00:00-08:00
    Supervisor: IA, with: 
    Log book#14 pages 22 - 22
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam from midnight to 3:30. Very high background. I 
    changed polarity in FS after a zero field run.
    2. Achievements :
      Data runs:
     7989 MRS 40 deg. Zero field run(50k); FS 12 deg. 1/4A(6k)
     7990 MRS 40 deg. 1050A (116k); FS 12 deg. 1/4A(22k)
     7991 MRS 40 deg. 1050A (86k); FS 12 deg. 1/4A(21k)
     7992 MRS 40 deg. 1050A (48k); FS 12 deg. Zero field (20k)
      Pedestal runs :
     7993 Non-TPC pedestal run
     7994 TPC pedestal run
    3. Mainframe L1458A turned off several times. I turn it on 
     without any problems.
       I changed polarity in FS from A to B.
       Low event rate for the DVM run (1-2 events/sec.).
       MRS angle is now 40.
    
    SITUATION:
      MRS: - need more 250k events on current settings (40 deg. 1050 A)
           - 500k events on 1050B 40 deg.
      FS:  - 200k events on current settings (12 deg. 1/4B)
           - 200k events on 12 deg. 1/4A
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Mon Feb 17 2003 - 08:06:06 EST